OpenStudy (anonymous):

what is the equation of the line that passes through (-2,-3) and is perpendicular to 2x-3y=6?

OpenStudy (mertsj):

The slope of 2x-3y=6 is 2/3 so the slope of the perpendicular is -3/2 Now write the equation of a line whose slope is -3/2 and contains the point (-2,-3)

OpenStudy (anonymous):

thank you..help me with another?

OpenStudy (anonymous):

write the equation of the line that passes through (-2,1) and is perpendicular to the line 3x-2y=5

OpenStudy (mertsj):

The slope of 3x-2y = 5 is 3/2 so the slope of the perpendicular is -2/3 So write the equation of a line whose slope is -2/3 and contains the point (-2,1)

OpenStudy (anonymous):

&& what is the equation of the line that passes through (-2,3) and is parallel to 2x+3y=6?

OpenStudy (mertsj):

if it is parallel, it has the same slope. The slope of 2x+3y = 6 is -2/3 so write the equation of line whose slope is -2/3 and contains (-2,3)